Circular No. 8367

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2004 J17–J19 AND K4–K12 (SOHO) Continuation to IAUC 8366 (DE = D. Evans; MM = M. Meyer; VB

= V. Bezugly):

Comet 2004 UT α_{2000} δ_{2000} Inst. F MPEC
C/2004 J17 May 15.435 3^{h}34.^{m}6 +19^{o}38′ C2  RK 2004-N05
C/2004 J18  15.561 3 35.7 +19 43 C2  RK 2004-N05
C/2004 J19  14.671 3 31.1 +17 29 C2  DE 2004-N16
C/2004 K4  18.410 3 45.5 +18 11 C2  RK 2004-N16
C/2004 K5  20.201 3 52.1 +18 33 C2  XZ 2004-N16
C/2004 K6  21.743 3 58.1 +18 48 C2  MB 2004-N16
C/2004 K7  23.088 4 03.2 +19 03 C2  DC 2004-N16
C/2004 K8  24.326 4 06.7 +19 11 C2  MM 2004-N16
C/2004 K9  24.826 4 09.2 +19 21 C2  VB 2004-N17
C/2004 K10  25.311 4 11.3 +19 26 C2  MM 2004-N17
C/2004 K11  27.726 4 20.8 +19 47 C2  XZ 2004-N17
C/2004 K12  30.446 4 33.0 +19 27 C3/2 RK 2004-N17

All are Kreutz comets except for C/2004 J17 and J18, which are Krachtgroup members. Corrigendum: On IAUC 8365, line 12, for 2004-D14. read 2004-M14.

POSSIBLE SUPERNOVA IN NGC 7377 Further to IAUC 8361, S. Park and W. Li report the LOSS discovery,

on KAIT images taken on June 28.43 and 29.41 UT, of a possible supernova (mag ∼ 17.8) located at α = 22^{h}47^{m}55^{s}.88, δ = −22^{o}19′40′′.5 (equinox 2000.0), which is 115′′.5 east and 57′′.3 south of the nucleus of NGC 7377. A KAIT image taken on 2003 Dec. 16.16 showed nothing at this position (limiting mag ∼ 19.0). S. Nakano, Sumoto, Japan, forwards the following position end figures from an unfiltered CCD image taken by K. Itagaki (Yamagata, Japan, 0.60-m f/5.7 reflector) on July 1.766: 55^{s}.88, 40′′.7 (mag 17.7). Itagaki adds that nothing is visible at this location on the Digitized Sky Survey or on his own earlier survey frames taken on 2002 Aug. 30 and 2003 Sept. 22 (limiting mag 18.5).

UNUSUAL CATACLYSMIC VARIABLE IN HERCULES Further to IAUC 8363, J. Bedient adds that the June 22 images were

taken at the 2.2-m reflector by J. Pittichova.

2004 July 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green